# Semitoric Families on Pentagon Spaces

### v1

@@ 5. The transition point $P$
We now show that when $t\in[0,t^{-})\cup(t^{+},1]$, both roots of $\chi_{1,1}(X)$ are negative by showing that $A(t)>0,B(t)>0$. To check that $B(t)>0$, it suffices to show that $g(t):=-r_{4}t^{2}+(r_{3}+r_{4}+r_{5})t+(r_{3}+2r_{4}-2r_{5})$ has no roots in $[0,1]$. Notice that $g(0)=r_{3}-r_{5}+2r_{4}-r_{5}>0$, $g(1)=2r_{3}+2r_{4}-r_{5}>0$ and $g$ describes a parabola opening downwards, so $g(t)$ has no roots in $[0,1]$, i.e. $B(t)=j^{2}g(t)^{2}>0$ for all $t\in[0,1]$. As for $A(t)$, we notice that the coefficient of $t^{2}$ is positive and the discriminant of $A(t)$ is
 $4j\Bigl((r_{3}-r_{5})((r_{3}+r_{5})^{2}+7r_{3}r_{4}+5r_{4}^{2})+(2r_{4}-r_{5})(5r_{3}r_{4}+3r_{4}^{2})+7r_{4}r_{5}(r_{5}-r_{4})+3r_{4}^{3}\Bigr)<0$
Hence, $A(t)$ is always positive. Consequently, for $t\in[0,t^{-})\cup(t^{+},1]$, the singular point $P$ is of elliptic-elliptic type, and for $t\in(t^{-},t^{+})$, $P$ is of focus-focus type. ∎
6. Proof of Theorem 1.1
We are now ready to prove our main theorem.
Proof of Theorem 1.1.
We write $F_{t}=(J,H_{t})$ and $M=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. By assumption, the moment images of $(J,\sqrt{H_{0}})$ and $(J,\sqrt{H_{1}})$ are both Delzant polygons. Hence, for $t=0,1$, the system is of toric type. Moreover, from the Delzant polygons (see Figure 1), $P$ is a maximum of $(H_{1})|_{J^{-1}(J(P))}$ and a minimum of $(H_{0})|_{J^{-1}(J(P))}$. Hence, it follows from Proposition 5.1 that $P$ is a transition point with transition time $t^{\pm}$.
We now show that for all $t\in[0,1]$, all other singular points of $F_{t}$ are non-degenerate of either elliptic-elliptic or elliptic-regular types. Fix a rank one point $m\in M$ that is not on the fixed surfaces $J^{-1}(|r_{1}-r_{2}|),J^{-1}(r_{1}+r_{2})$. By Lemma 2.5, the $S^{1}$-action generated by $J$ acts freely on $m$. By Proposition 4.2, if $c=J(m)\neq r_{3}+r_{4}-r_{5}$, then $[m]$ is an elliptic (in the Morse sense) singular point of $H_{t}^{\textrm{red},c}$ on $M^{\textrm{red},c}:=J^{-1}(c)/S^{1}$. Hence, by Lemma 3.7, $m$ is an elliptic-regular singularity of $F_{t}$. If $c=r_{3}+r_{4}-r_{5}$, then $M^{\textrm{red},c}$ is not a manifold. However, since $S^{1}$ acts freely on $m$, there is an $S^{1}$-invariant neighborhood around the orbit of $m$ on which $S^{1}$ acts freely. Hence, one can do symplectic reduction locally, and verify that $[m]$ is an elliptic singularity of $H_{t}^{\textrm{red},c}$ by computing its Hessian in the same way as in the proof of Proposition 4.2. Therefore, any rank one points that are not on the fixed surfaces are of elliptic-regular type. Finally, Proposition 3.9 implies that any rank one point on the fixed surfaces is of elliptic-regular type, and Proposition 3.9 and Proposition 4.2 together imply that any rank zero point on the fixed surfaces is of elliptic-elliptic type. Therefore, we conclude that for $t\in[0,1]\setminus\{t^{-},t^{+}\}$, $(M,\omega,F_{t})$ is a semitoric system, and when $t=t^{\pm}$, there are no degenerate singular points in $M\setminus\{P\}$. ∎

### v2

@@ 5. The transition points
We now show that when $t\in[0,t_{P}^{-})\cup(t_{P}^{+},1]$, both roots of $\chi_{\lambda,1}(X)$ are negative by showing that $A(t)>0,B(t)>0$. To check that $B(t)>0$, it suffices to show that $g(t):=-r_{4}t^{2}+r_{3}\lambda t+r_{4}t+r_{5}\lambda t+j_{P}\lambda^{2}+r_{4}\lambda-r_{5}\lambda$ has no roots in $[0,1]$. Since $\lambda>\frac{r_{5}-r_{4}}{j_{P}}>0$, $g(0)=\lambda(j_{P}\lambda+r_{4}-r_{5})$ and $g(1)=\lambda(j_{P}\lambda+r_{3}+r_{4})$ are both positive. Since $g$ describes a parabola opening downward, $g(t)$ has no roots in $[0,1]$, i.e., $B(t)=j_{P}^{2}g(t)^{2}>0$ for all $t\in[0,1]$. As for $A(t)$, we notice that the coefficient of $t^{2}$ is positive and the discriminant $\Delta$ of $A(t)$ is
$\Delta=-\,4j_{P}\Bigl([j_{P}^{3}+2r_{4}j_{P}^{2}+(r_{4}^{2}+4r_{3}r_{5})j_{P}]\lambda^{2}+2r_{4}j_{P}(j_{P}+r_{4})\lambda-2r_{3}r_{4}r_{5}+r_{4}^{2}j_{P}\Bigr).$
Since $\lambda>\frac{r_{3}r_{5}}{j_{P}(j_{P}+r_{4})}$, we have $2r_{4}j_{P}(j_{P}+r_{4})\lambda-2r_{3}r_{4}r_{5}>0$, so $\Delta$ is negative and $A(t)$ is always positive. Consequently, $P$ is of elliptic–elliptic type for $t\in[0,t_{P}^{-})\cup(t_{P}^{+},1]$, and of focus–focus type for $t\in(t_{P}^{-},t_{P}^{+})$. This concludes the proof of case (1).
To derive the proof of case (2) from case (1), we first notice that the inequalities in case (Q) imply that $r_{3}>r_{4}$ and $j_{Q}>0$. Then one replaces $P$ with $Q$, $r_{3}$ with $r_{3}$, and $r_{5}$ with $r_{5}$ in the preceding Hessian computations and sign analysis, and argues in the same manner that the relevant quantities remain positive or negative under the choice $\lambda>\max(\frac{r_{3}-r_{4}}{j_{Q}},\frac{r_{3}r_{5}}{j_{Q}(r_{4}+j_{Q})})$. Finally, the last case can be proved by noticing that the inequalities in case (P+Q) imply that $r_{3},r_{5}>r_{4}$, $j_{P},j_{Q}>0$ and joining the proof of case (1) and case (2). ∎